OpenStudy (anonymous):

Solve: negative 1 over 2 x + 3 = –x + 7 1 over 2 8 –8 2

OpenStudy (anonymous):

are we solving for x?

OpenStudy (anonymous):

@mz.inno143

OpenStudy (michele_laino):

hint: if I multiply both sides of your equation by 2, I get this: \[\Large 2\left( { - \frac{1}{2}x + 3} \right) = 2\left( { - x + 7} \right)\]

OpenStudy (anonymous):

answer is 8/3

OpenStudy (anonymous):

the link shows you the steps to your problem and its also like an online problem solver, i tend to use it a lot and find it very helpful :)

OpenStudy (michele_laino):

the equation is: negative 1 over 2... @LaurenRoberts

OpenStudy (anonymous):

well then @mz.inno143 could just use the link i gave cause it takes you to an online math problem solver and gives you steps :)

OpenStudy (anonymous):

thank you guys very great help